# Asymptotically tight Lagrangian dual of smooth nonconvex problems via improved error bound of Shapley-Folkman Lemma

### v1

@@ 4.3 Improved error bounds via local geometry
$\leq 1+\frac{L}{2}\sum_{i\in\mathcal{I}}\Xi^{2}(F_{i}).$
This implies there exists some point in $\sum_{i\in\mathcal{I}}A_{i}\cap F_{i}$ whose distance defined by $\left\lVert\cdot\right\rVert_{\mathcal{H}}$ to $\mathbf{p}$ is at most $\sqrt{1+\frac{L}{2}\sum\limits_{i\in\mathcal{I}}\Xi^{2}(F_{i})}$. Therefore, by the positive homogeneousness of $\left\lVert\cdot\right\rVert_{\mathcal{H}}$, the $\left\lVert\cdot\right\rVert_{\mathcal{H}}$ distance of $\mathbf{p}$ to $\left(\sum_{i\in\mathcal{I}}A_{i}\cap F_{i}\right)+\mathcal{H}$ is at most $\sqrt{1+\frac{L}{2}\sum\limits_{i\in\mathcal{I}}\Xi^{2}(F_{i})}-1$.
@@
Let $A_{1},A_{2},\dots,A_{k}$ be nonempty subsets of $\mathbb{R}^{n}$ and $\Xi(A_{i})\leq\beta,\forall i\in[k]$ such that ${\mathbf{conv}}(A_{i})$ is pointed and closed. If every extreme point of $A_{i}$ has a $l_{2}$ ball with radius $r_{i}$ ($r_{i}$ can be zero) as hidden convex component, let $r_{*}=\inf_{\mathcal{I}\subseteq[k]:|\mathcal{I}|=n}\sum_{i\in[k]\setminus\mathcal{I}}r_{i}$. Then it follows that
@@
Proof 4.20.
To bound $\Phi\left(\sum_{i=1}^{k}A_{i}\right)$, it suffices to show that for any $\mathbf{p}\in\sum_{i=1}^{k}{\mathbf{conv}}(A_{i})$, we have $\Phi\left(\mathbf{p},\sum_{i=1}^{k}A_{i}\right)\leq\sqrt{r_{*}^{2}+\dfrac{1}{2}n\beta^{2}}-r_{*}$. By Lemma 4.6, there exists face $F_{i}$ of ${\mathbf{conv}}(A_{i})$ such that $\mathbf{p}\subseteq\sum_{i=1}^{k}F_{i}$ and $\sum_{i=1}^{k}\dim(F_{i})=n$. By the hypothesis, we can find $H_{i}=\mathcal{B}(\textbf{0},r_{i})$ as hidden convex component associated with $F_{i}$ for all $i\in[k]\setminus\mathcal{I}$ such that $\mathcal{H}=\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}=\mathcal{B}(\textbf{0},r_{*})$. Then $\left\lVert\mathbf{x}\right\rVert_{\mathcal{H}}=\dfrac{1}{r_{*}}\left\lVert\mathbf{x}\right\rVert$ and $\left\lVert\cdot\right\rVert_{\mathcal{H}}^{2}$ is $\dfrac{1}{r_{*}^{2}}$-smooth. Lemma 4.17 implies that
$\Phi^{\mathcal{H}}\left(\sum_{i=1}^{k}F_{i},\sum_{i=1}^{k}A_{i}\right)\leq\sqrt{1+\dfrac{1}{2r_{*}^{2}}\sum_{i\in\mathcal{I}}^{n}\Xi^{2}(F_{i})}-1.$
@@
$\Phi\left(\sum_{i=1}^{k}F_{i},\sum_{i=1}^{k}A_{i}\right)$ $\leq\sqrt{r_{*}^{2}+\dfrac{1}{2}\sum_{i\in\mathcal{I}}^{n}\Xi^{2}(F_{i})}-r_{*}$
$\leq\sqrt{r_{*}^{2}+\dfrac{1}{2}\sum_{i\in\mathcal{I}}^{n}\beta^{2}}-r_{*}$
where the last inequality uses Claim 2. Since $\mathbf{p}$ is arbitrary, this implies that $\Phi\left(\sum_{i=1}^{k}A_{i}\right)\leq\sqrt{r_{*}^{2}+\dfrac{1}{2}n\beta^{2}}-r_{*}.$. The rest of statement directly follows that $r_{*}\leq(k-n)r$.
We end the section with serval examples and remarks.
@@
It requires the existence of smooth hidden convex components only around the extreme points of each nonconvex set in order to ensure that their Minkowski sums converge to a convex set. Such smooth hidden components need not exist around non-extreme points. To illustrate this, consider the $\mathbb{R}^{2}$ examples $T_{1}$ and $T_{2}$ shown in Figures 8 and 8. Both sets contain some points on which no smooth hidden convex components are associated with. Nevertheless, one can verify that $\sum_{i=1}^{k}T_{1}$ converges to a convex set as $k\to\infty$, whereas $\sum_{i=1}^{k}T_{2}$ does not. The difference arises from the behavior at extreme points: every extreme point of ${\mathbf{conv}}(T_{1})$ admits a smooth hidden convex component (e.g., a ball contained in $T_{1}$ in a neighborhood of that point), while ${\mathbf{conv}}(T_{2})$ possesses five extreme points for which no such smooth hidden convex component exists.
@@ 5 Asymptotically tight Lagrangian dual of smooth nonconvex problems
In this section, we aim to leverage Theorem 4 and Lemma 4.17 to establish the asymptotic tightness of problem (1). The key insight is that the smooth hidden convex components naturally emerge at the extreme points of the convex hull of the epigraphs of smooth functions. Under mild assumptions on the constraint matrices $B^{(i)}$, the smoothness of these hidden convex components is approximately preserved under projection, thereby implying that $\Phi(\mathcal{P})$ vanishes as $k$ increases. We rewrite (4):
@@
Note that we can simply choose $s_{i}=n$ in (7) to model the block-structured smooth nonconvex problem without sparsity constrained. On the other hand, $f^{(i)}$ is a function from $\mathbb{R}^{n}$ to $\mathbb{R}$ and therefore is proper.
@@
A function $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called 1-coercive if
$\frac{f(x)}{\|x\|}\;\longrightarrow\;+\infty\qquad\text{as }\|x\|\to\infty.$
@@
1-coercive function is commonly used in analysis and optimization. Standard example includes strictly convex quadratic functions and univariate polynomial with a positive leading coefficient. The existence of asymptotically tight Lagrangian dual requires two additional assumption. Assumption 4 requires the nonconvexity of each block in (4) is bounded. Assumption 5 implies that ${\mathbf{conv}}(\mathcal{X}^{(i)}_{\textbf{s}})$ is closed and pointed in Lemma 5.25, which ensures Assumption 2 and Assumption 3.
@@
Each $f^{(i)}(\cdot)$ is $1$-coercive and closed.
@@
Under assumption 5, ${\mathbf{conv}}(\mathcal{X}^{(i)}_{\textbf{s}})$ is closed and pointed.
Proof 5.26.
@@
Let $\chi_{s_{i}}(\mathbf{x}):=\begin{cases}\infty&\text{ if }\left\lVert\mathbf{x}^{(i)}\right\rVert_{0}\geq s_{i}\\ 0&\text{ if otherwise}\end{cases}.$ Let $\tilde{f}^{(i)}(\mathbf{x}):=f^{(i)}(\mathbf{x})+\chi_{s_{i}}(\mathbf{x})$. By the definition of $1$-coercive, it is straightforward that $\tilde{f}^{(i)}(\mathbf{x})$ is $1$-coercive and closed. Moreover, it is clear that $\operatorname{\operatorname{epi}}\tilde{f}^{(i)}(\mathbf{x})=\mathcal{X}^{(i)}_{\textbf{s}}$. It is well known that the convex hull of epigraph of proper, $1$-coercive and closed function is closed and therefore ${\mathbf{conv}}\left(\mathcal{X}^{(i)}_{\textbf{s}}\right)$ is closed [hiriart1993convex]. Suppose that ${\mathbf{conv}}\left(\mathcal{X}^{(i)}_{\textbf{s}}\right)$ is not pointed, we seek contradiction. We first observe that the definition of $1$-coercive implies that there exists some $lb$ such that $lb\leq\tilde{f}^{(i)}(\mathbf{x})$. This implies that $t\geq lb,\forall(t,\mathbf{x})\in{\mathbf{conv}}(\mathcal{X}_{\textbf{s}}^{(i)})$. Therefore, any line contained in ${\mathbf{conv}}(\mathcal{X}_{\textbf{s}}^{(i)})$ must take form of $l(t):=(t_{0},\mathbf{x}_{0})+t(0,\mathbf{r})\forall t\in\mathbb{R}^{n}$ for some nonzero $\mathbf{r}$. Since $\tilde{f}^{(i)}(\mathbf{x})$ is $1$-coercive, there exists some $R>1$ such that $\left\lVert\mathbf{x}\right\rVert_{2}\geq R\implies\tilde{f}^{(i)}(\mathbf{x})\geq(|lb|+|t_{0}|+1)\left\lVert\mathbf{x}\right\rVert$. Choose $t$ sufficiently large so that $\left\lVert\mathbf{x}_{0}+t\mathbf{r}\right\rVert_{2}\geq(n+2)R$, since $(t_{0},\mathbf{x}_{0}+t\mathbf{r})\in{\mathbf{conv}}(\mathcal{X}_{\textbf{s}}^{(i)})$, this implies that $\mathbf{x}_{0}+t\mathbf{r}=\sum_{i\in\mathcal{K}}\lambda_{i}(t_{i},\mathbf{x}_{i})$ for some $(t_{i},\mathbf{x}_{i})\in\mathcal{X}_{\textbf{s}}^{(i)}$ and $\lambda$ from a simplex of dimension $|\mathcal{K}|$. Since $\left\lVert\cdot\right\rVert_{2}$ is subadditive and positive homogeneous, this implies that $\left\lVert\mathbf{x}_{0}+t\mathbf{r}\right\rVert\leq\sum_{i\in\mathcal{K}}\lambda_{i}\left\lVert\mathbf{x}_{i}\right\rVert$. By Carathéodory’s Theorem, we have $|\mathcal{K}|\leq(n+2)$. Since $\left\lVert\mathbf{x}_{0}+t\mathbf{r}\right\rVert\geq(n+2)R$, a simple pigeonhole argument implies that there exists some $j\in\mathcal{K}$, such that $\lambda_{j}\left\lVert\mathbf{x}_{j}\right\rVert\geq R$. Finally, this show that
@@
$\left\lVert\mathbf{x}_{j}\right\rVert\geq R\implies\lambda_{j}t_{j}\geq\dfrac{R}{\left\lVert\mathbf{x}\right\rVert}t_{j}\geq R(|lb|+|t_{0}|+1)\geq|lb|+|t_{0}|+1.$
@@
$\sum_{i\in\mathcal{K}}\lambda_{i}t_{i}=\lambda_{j}t_{j}+\sum_{i\in\mathcal{K}\setminus\{j\}}\lambda_{i}t_{i}\geq|lb|+|t_{0}|+1+lb>t_{0},$
@@
The last ingredient for applying Lemma 4.17 is to show that there exists a smooth hidden convex component associated with each extreme point of ${\mathbf{conv}}\bigl(\mathcal{P}_{\mathbf{s}}^{(i)}\bigr)$. We obtain this from the following observation. The set $\mathcal{P}_{\mathbf{s}}^{(i)}$ can be viewed as a linear projection of a union of epigraphs of certain smooth functions. By the definition of smoothness, for any point in such an epigraph we can find a ball that lies entirely in the epigraph and contains this point by Claim 3. Under the linear projection induced by the coupling matrix $B^{(i)}$, this ball is mapped to an ellipsoid in $\mathcal{P}_{\mathbf{s}}^{(i)}$, which serves as the smooth hidden convex component we are looking for. The maximal distortion under this linear projection is quantified by $\mathcal{L}_{s}(B^{(i)})$, defined below.
@@
Given a matrix $B\in\mathbb{R}^{m\times n}$ and let $\tilde{B}\in\mathbb{R}^{(m+1)\times n}:=\begin{bmatrix}\textbf{1}^{\top}\\ B\end{bmatrix}$ be a matrix with appending a row of all ones. We define the projection factor with respect to sparsity level $s$ as
$\mathcal{L}_{s}(B):=\inf_{S\subseteq[n]:|S|=s}\sigma_{\inf}\left(\tilde{B}_{S}(\tilde{B}_{S})^{\top}\right)$
where $\sigma_{\inf}(\cdot)$ is the minimal singular value and $\tilde{B}_{S}\in\mathbb{R}^{(m+1)\times|S|}$ is the submatrix of $\tilde{B}$ that consists columns of $B$ with indices in $S$.
@@
To obtain a nondegenerate hidden convex component, we require $\mathcal{L}_{s}(B^{(i)})>0$. This condition is equivalent to the existence of $m+1$ affinely independent vectors among every collection of $s^{(i)}$ columns of $B^{(i)}$. A necessary condition for this is $s^{(i)}\geq m+1$. Apart from this simple dimensional requirement, the condition $\mathcal{L}_{s}(B^{(i)})>0$ is not very restrictive: it holds generically and can be ensured almost surely by adding an arbitrarily small smooth random perturbation to $B^{(i)}$ whenever $s^{(i)}\geq m+1$.

### v2

@@ 4.3 Improved error bounds via local geometry
$\leq 1+\frac{L}{2}\sum_{i\in\mathcal{I}}\Xi^{2}(A_{i}\cap F_{i}).$
This implies there exists some point in $\sum_{i\in\mathcal{I}}A_{i}\cap F_{i}$ whose distance defined by $\left\lVert\cdot\right\rVert_{\mathcal{H}}$ to $\mathbf{p}$ is at most $\sqrt{1+\frac{L}{2}\sum\limits_{i\in\mathcal{I}}\Xi^{2}(A_{i}\cap F_{i})}$. Therefore, by the positive homogeneousness of $\left\lVert\cdot\right\rVert_{\mathcal{H}}$, the $\left\lVert\cdot\right\rVert_{\mathcal{H}}$ distance of $\mathbf{p}$ to $\left(\sum_{i\in\mathcal{I}}A_{i}\cap F_{i}\right)+\mathcal{H}$ is at most $\sqrt{1+\frac{L}{2}\sum\limits_{i\in\mathcal{I}}\Xi^{2}(A_{i}\cap F_{i})}-1$.
@@
Let $A_{1},A_{2},\dots,A_{k}$ be nonempty subsets of $\mathbb{R}^{n}$ and $\Xi(A_{i})\leq\beta$, for all $i\in[k]$ such that ${\mathbf{conv}}(A_{i})$ is closed and pointed. Suppose every extreme point of $A_{i}$ has an associated $l_{2}$–ball with radius $r_{i}$ ($r_{i}$ can be zero) as a hidden convex component. Let $r_{*}=\min_{\mathcal{I}\subseteq[k]:|\mathcal{I}|=n}\sum_{i\in[k]\setminus\mathcal{I}}r_{i}$. Then it follows that
@@
Proof 4.23.
To bound $\Phi\left(\sum_{i=1}^{k}A_{i}\right)$, it suffices to show that for any $\mathbf{p}\in\sum_{i=1}^{k}{\mathbf{conv}}(A_{i})$, we have $\Phi\left(\mathbf{p},\sum_{i=1}^{k}A_{i}\right)\leq\sqrt{r_{*}^{2}+\dfrac{1}{2}n\beta^{2}}-r_{*}$. By Lemma 4.7, there exists face $F_{i}$ of ${\mathbf{conv}}(A_{i})$ such that $\mathbf{p}\in\sum_{i=1}^{k}F_{i}$ and $\sum_{i=1}^{k}\dim(F_{i})\leq n$.
By the hypothesis, after suitable translation of all the sets, we can find $H_{i}=\mathcal{B}(\textbf{0},r_{i})$ as hidden convex component associated with $F_{i}$ for all $i\in[k]\setminus\mathcal{I}$ such that $\mathcal{H}=\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}=\mathcal{B}(\textbf{0},r_{*})$.
Then $\left\lVert\mathbf{x}\right\rVert_{\mathcal{H}}=\dfrac{1}{r_{*}}\left\lVert\mathbf{x}\right\rVert$ and $\left\lVert\cdot\right\rVert_{\mathcal{H}}^{2}$ is $\dfrac{1}{r_{*}^{2}}$-smooth. Lemma 4.20 implies that
$\Phi^{\mathcal{H}}\left(\sum_{i=1}^{k}F_{i},\sum_{i=1}^{k}A_{i}\right)\leq\sqrt{1+\dfrac{1}{2r_{*}^{2}}\sum_{i\in\mathcal{I}}\Xi^{2}(A_{i}\cap F_{i})}-1.$
@@
$\Phi\left(\sum_{i=1}^{k}F_{i},\sum_{i=1}^{k}A_{i}\right)$ $\leq\sqrt{r_{*}^{2}+\dfrac{1}{2}\sum_{i\in\mathcal{I}}\Xi^{2}(A_{i}\cap F_{i})}-r_{*}$
$\leq\sqrt{r_{*}^{2}+\dfrac{1}{2}\sum_{i\in\mathcal{I}}\beta^{2}}-r_{*},$
where the last inequality uses Claim 4. Since $\mathbf{p}$ is arbitrary, this implies that $\Phi\left(\sum_{i=1}^{k}A_{i}\right)\leq\sqrt{r_{*}^{2}+\dfrac{1}{2}n\beta^{2}}-r_{*}$. The rest of statement directly follows that $r_{*}\geq(k-n)r$.
We end the section with several examples and remarks.
@@
To ensure that the Minkowski sums converge to a convex set, it is sufficient that smooth, hidden convex components exist only in neighborhoods of the extreme points of convex hull of each nonconvex set. Such components need not exist around non-extreme points of the convex hulls. To illustrate this, consider the $\mathbb{R}^{2}$ examples $T_{1}$ and $T_{2}$ shown in Figures 8,  8,  8 and 8. Both sets contain some points on which no smooth hidden convex components are associated with. Nevertheless, one can verify that $\sum_{i=1}^{k}T_{1}$ converges to a convex set as $k\to\infty$, whereas $\sum_{i=1}^{k}T_{2}$ does not. The difference arises from the behavior at extreme points: every extreme point of ${\mathbf{conv}}(T_{1})$ admits a smooth hidden convex component (e.g., a ball contained in $T_{1}$ in a neighborhood of that point), while ${\mathbf{conv}}(T_{2})$ possesses five extreme points for which no such smooth hidden convex component exists.
@@ 5 Asymptotically tight Lagrangian dual of smooth nonconvex problems
In this section, we aim to leverage Theorem 4 and Lemma 4.20 to establish the asymptotic tightness of problem (4). The key insight is that the smooth hidden convex components naturally emerge at the extreme points of the convex hull of the epigraphs of smooth functions. Under mild assumptions on the constraint matrices $B^{(i)}$, the smoothness of these hidden convex components is approximately preserved under projection, thereby implying that $\Phi(\mathcal{P})$ vanishes as $k$ increases. We rewrite (4) as:
@@
Note that we can simply choose $s_{i}=n$ in (8) to model the block-structured smooth nonconvex problem without sparsity constraints.
@@
[21] A function $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called 1-coercive if
$\frac{f(x)}{\|x\|_{2}}\;\longrightarrow\;+\infty\qquad\text{as }\|x\|_{2}\to\infty.$
@@
The notion of 1-coercive functions is commonly used in analysis and optimization [24]. Some example includes strictly convex quadratic functions and univariate polynomials of even degree with a positive leading coefficient.
The existence of an asymptotically tight Lagrangian dual requires two additional assumptions. Assumption 3 requires the nonconvexity of each block in (4) is bounded. Assumption 4 implies that ${\mathbf{conv}}\left(\mathcal{P}^{(i)}_{\textbf{s}}\right)$ is closed and shown in Lemma 5.31.
@@
A function $f(\mathbf{x}):\mathbb{R}^{n}\to(-\infty,+\infty]$ is called proper if there exists some $\mathbf{x}_{0}$ such that $f(\mathbf{x}_{0})<\infty$.
@@
Each $f^{(i)}(\cdot)$ is $1$-coercive, closed and proper and $f^{(i)}(\textbf{0})<\infty$.
@@
Under Assumption 4, ${\mathbf{conv}}\left(\mathcal{X}^{(i)}_{\textbf{s}}\right)$ is closed and pointed. Moreover, $\text{rec}\left({\mathbf{conv}}\left(\mathcal{X}_{\textbf{s}}^{(i)}\right)\right)=\{(t,\text{0}):t\geq 0\}$.
Proof 5.29.
@@
Let $\chi_{s_{i}}(\mathbf{x}):=\begin{cases}\infty&\text{ if }\left\lVert\mathbf{x}^{(i)}\right\rVert_{0}\geq s_{i}{+1}\\ 0&\text{ if otherwise}\end{cases}.$ Let $\tilde{f}^{(i)}(\mathbf{x}):=f^{(i)}(\mathbf{x})+\chi_{s_{i}}(\mathbf{x})$. By the definition of $1$-coercive, it is straightforward that $\tilde{f}^{(i)}(\mathbf{x})$ is $1$-coercive and closed. Since $f^{(i)}(\textbf{0})<\infty$, $\tilde{f}^{(i)}(\mathbf{x})$ is also proper. Moreover, it is clear that $\operatorname{\operatorname{epi}}\tilde{f}^{(i)}(\mathbf{x})=\mathcal{X}^{(i)}_{\textbf{s}}$. It is well known that the convex hull of the epigraph of a proper, $1$-coercive and closed function is closed and therefore ${\mathbf{conv}}\left(\mathcal{X}^{(i)}_{\textbf{s}}\right)$ is closed [18]. Suppose, for the sake of contradiction, that ${\mathbf{conv}}\left(\mathcal{X}^{(i)}_{\textbf{s}}\right)$ is not pointed. We first observe that the definition of $1$-coercive-ness implies that there exists some $lb$ such that $lb\leq\tilde{f}^{(i)}(\mathbf{x})$. This implies that $t\geq lb$, for all $(t,\mathbf{x})\in{\mathbf{conv}}\left(\mathcal{X}_{\textbf{s}}^{(i)}\right)$. Therefore, any line contained in ${\mathbf{conv}}\left(\mathcal{X}_{\textbf{s}}^{(i)}\right)$ must take the form of $l(\gamma):=(t_{0},\mathbf{x}_{0})+\gamma(0,\mathbf{r}),\ \gamma\in\mathbb{R}^{n}$ for some nonzero $\mathbf{r}$. To also establish the properties of $\text{rec}\left({\mathbf{conv}}\left(\mathcal{X}_{\textbf{s}}^{(i)}\right)\right)$, we show the stronger statement that ${\mathbf{conv}}\left(\mathcal{X}_{\textbf{s}}^{(i)}\right)$ does not contain any ray taking the form of $l^{+}(\gamma):=(t_{0},\mathbf{x}_{0})+\gamma(0,\mathbf{r}),\ \gamma\in\mathbb{R}_{+}$ for some nonzero $\mathbf{r}$. This implies that ${\mathbf{conv}}\left(\mathcal{X}_{\textbf{s}}^{(i)}\right)$ is pointed. Combined with the fact that the function is also lower bounded by some constant, this also implies that $\text{rec}\left({\mathbf{conv}}\left(\mathcal{X}_{\textbf{s}}^{(i)}\right)\right)=\{(t,\text{0}):t\geq 0\}$. The argument is as follows.
Suppose such ray $l^{+}(\gamma)$ exists, since $\tilde{f}^{(i)}(\mathbf{x})$ is $1$-coercive, there exists some $R>1$ such that $\left\lVert\mathbf{x}\right\rVert_{2}\geq R\implies\tilde{f}^{(i)}(\mathbf{x})\geq(|lb|+|t_{0}|+1)\left\lVert\mathbf{x}\right\rVert_{2}$. Choose $|\gamma|$ sufficiently large so that $\left\lVert\mathbf{x}_{0}+\gamma\mathbf{r}\right\rVert_{2}\geq(n+2)R$, since $(t_{0},\mathbf{x}_{0}+\gamma\mathbf{r})\in{\mathbf{conv}}\left(\mathcal{X}_{\textbf{s}}^{(i)}\right)$, this implies that $\mathbf{x}_{0}+\gamma\mathbf{r}=\sum_{j\in[n+2]}\lambda_{j}(t_{j},\mathbf{x}_{j})$ for some $(t_{j},\mathbf{x}_{j})\in\mathcal{X}_{\textbf{s}}^{(i)}$ and $\lambda$ belongs to a simplex of dimension $n+2$. Since $\left\lVert\cdot\right\rVert_{2}$ is subadditive and positively homogeneous, this implies that $\left\lVert\mathbf{x}_{0}+\gamma\mathbf{r}\right\rVert_{2}\leq\sum_{i\in[n+2]}\lambda_{j}\left\lVert\mathbf{x}_{j}\right\rVert_{2}$. Since $\left\lVert\mathbf{x}_{0}+\gamma\mathbf{r}\right\rVert_{2}\geq(n+2)R$, a simple pigeonhole argument implies that there exists some $j_{*}\in[n]$, such that $\lambda_{j_{*}}\left\lVert\mathbf{x}_{j_{*}}\right\rVert_{2}\geq R$. Since $\lambda_{j_{*}}\in[0,1]$, this further shows that $\left\lVert\mathbf{x}_{j_{*}}\right\rVert_{2}\geq R$ and by the choice of $R$, this implies that $t_{j_{*}}\geq(|lb|+|t_{0}|+1)\left\lVert\mathbf{x}_{j_{*}}\right\rVert_{2}$. This shows that
@@
$\lambda_{j_{*}}\left\lVert\mathbf{x}_{j_{*}}\right\rVert_{2}\geq R\implies\lambda_{j_{*}}t_{j_{*}}\geq\dfrac{R}{\left\lVert\mathbf{x}_{j_{*}}\right\rVert_{2}}t_{j_{*}}\geq R(|lb|+|t_{0}|+1)\geq|lb|+|t_{0}|+1.$
@@
$\sum_{j\in[n+2]}\lambda_{j}t_{j}=\lambda_{j_{*}}t_{j_{*}}+\sum_{j\in[n+2]\setminus\{{j_{*}}\}}\lambda_{j}t_{j}\geq|lb|+|t_{0}|+1+(1-\lambda_{j_{*}})lb>t_{0},$
@@
Lemma 5.30 (Theorem 9.1 in [26]).
Let $C$ be a closed pointed convex set of $\mathbb{R}^{n}$. For any linear map $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$, if $\text{rec}(C)\cap\ker(\mathcal{A})=\{0\}$, then $\mathcal{A}(C)$ is closed and pointed.
Lemma 5.31.
Under assumption 4, ${\mathbf{conv}}\left(\mathcal{P}^{(i)}_{\textbf{s}}\right)$ is closed and pointed.
Proof 5.32.
By Claim 3, ${\mathbf{conv}}\left(\mathcal{P}^{(i)}_{\textbf{s}}\right)$ can be viewed as the image of ${\mathbf{conv}}\left(\mathcal{X}^{(i)}_{\textbf{s}}\right)$ under the following linear map $\mathcal{A}$: $(t,\mathbf{x})\to\begin{bmatrix}1&\\ &B\end{bmatrix}(t,\mathbf{x})$. In this case, we have $\text{rec}\left({\mathbf{conv}}\left(\mathcal{X}^{(i)}_{\textbf{s}}\right)\right)\cap\ker(\mathcal{A})=\{\textbf{0}\}$ because any non-zero $\mathbf{v}\in\text{rec}\left({\mathbf{conv}}\left(\mathcal{X}^{(i)}_{\textbf{s}}\right)\right)$ has form of $(r,\textbf{0})$ for some $r>0$ by Lemma 5.28 and it is clear that $\mathcal{A}\mathbf{v}\neq\textbf{0}$ in this case. Applying Lemma 5.30 and Lemma 5.28 yields the desired statement.
The last ingredient for applying Theorem 4.22 is to show that there exists a smooth hidden convex component ($l_{2}$-ball) associated with each extreme point of ${\mathbf{conv}}\bigl(\mathcal{P}_{\mathbf{s}}^{(i)}\bigr)$. We obtain this from the following observation. The set $\mathcal{P}_{\mathbf{s}}^{(i)}$ can be viewed as a linear projection of a union of epigraphs of certain smooth functions. By the definition of smoothness, for any point in such an epigraph we can find a ball that lies entirely in the epigraph and contains this point, as established in Claim 5 below. Under the linear projection induced by the coupling matrix $B^{(i)}$, this ball is mapped to an ellipsoid in $\mathcal{P}_{\mathbf{s}}^{(i)}$. This ellipsoid provides a desired smooth hidden convex component. We then extract an $l_{2}$-ball contained within this ellipsoid, which ultimately serves as the hidden convex component used in our proof. The maximal distortion under this linear projection and the choice $l_{2}$-ball contained in this ellipsoid is quantified by $\mathcal{L}_{s}(B^{(i)})$, defined below.
@@
Given a matrix $B\in\mathbb{R}^{m\times n}$ and let $\tilde{B}_{S}\in\mathbb{R}^{(m+1)\times(|S|+1)}:=\begin{bmatrix}1&\textbf{0}\\ \textbf{0}&B_{S}\end{bmatrix}$ be a matrix where $B_{S}$ is the submatrix of $B$ that consists of columns of $B$ with indices in $S$. We define the projection factor with respect to sparsity level $s$ as
$\mathcal{L}_{s}(B):=\inf_{S\subseteq[n]:|S|=s}\sigma_{\diamond}\left(\tilde{B}_{S}(\tilde{B}_{S})^{\top}\right)$
where $\sigma_{\diamond}(A):=\begin{cases}0&\text{ if }A\text{ is not positive definite};\\ \frac{\sqrt{\sigma_{\min}(A^{-1/2})}}{\sigma_{\max}(A^{-1/2})}&\text{ otherwise};\end{cases}$ and $\sigma_{\min}(\cdot),\sigma_{\max}(\cdot)$ are the minimal singular value and maximal singular value, respectively. $A^{1/2}$ denotes the principal square root of a positive definite matrix $A$.
@@
To obtain a nondegenerate hidden convex component, we require $\mathcal{L}_{s}(B^{(i)})>0$. This condition is equivalent to the existence of $m$ linear independent columns among every possible collection of $s^{(i)}$ columns of $B^{(i)}$. A necessary condition for this is $s^{(i)}\geq m$. Apart from this simple dimensional requirement, the condition $\mathcal{L}_{s}(B^{(i)})>0$ is not very restrictive: it holds generically and can be ensured almost surely by adding an arbitrarily small smooth random perturbation to $B^{(i)}$ whenever $s^{(i)}\geq m$.